OpenStudy (anonymous):

if P and Q is 12 cm and P and S 5 cm what would be the value of S and R

OpenStudy (triciaal):

do you have more information? what is the shape of this figure or the relationship between S and R

OpenStudy (triciaal):

P + Q = 12 P + S = 5 2 P + Q + S = 17 Q - S = 7 ; S = Q - 7 S = 17 -(2 P + Q) = 17 - 2 P - Q Q - 7 = 17 - 2 P - Q 2 Q = 24 - 2 P Q = 12 - P P - Q = -2 so S = Q - 7 and no idea where R fits in